# The NHIM development scenario in an asymmetric binary system of dwarf galaxies

### v1

@@ Abstract
We study the development scenario of a three-degree-of-freedom Hamiltonian system. We observe how the fundamental normally hyperbolic invariant manifolds (NHIMs) form a homoclinic/heteroclinic tangle. As the energy increases, the NHIMs begin to lose normal hyperbolicity and decay, creating tangential transient effects. We also observe a certain kind of coordination between the development scenarios of these NHIMs. We analyse this phenomenon using Poincaré maps and phase space indicator functions. As an example of demonstration, we use a relatively simple model of a test particle moving in the gravitational field of a pair of interacting dwarf galaxies.
@@ 1 Introduction
In dynamical systems showing chaos, unstable invariant subsets are the geometrical structures in the phase space which determine the global dynamics and, in particular, the creation of chaos to a large extent. As the simplest example, we can think of the unstable fixed points in a 2-dimensional map whose stable and unstable manifolds build up the chaotic invariant set, which can be displayed pictorially as a horseshoe construction traced out by the intersecting stable and unstable manifolds of the most important unstable fixed points. In maps with a domain of higher dimension, the role played in 2-dimensional maps by unstable fixed points is taken over by normally hyperbolic invariant manifolds (standard abbreviation NHIMs) of codimension 2; for general information on NHIMs see [1]. These subsets have stable and unstable manifolds of codimension 1, which can form walls, channels, and tubes in the domain of the map and thereby direct the global dynamics. This is the direct generalisation of the role of stable and unstable manifolds of hyperbolic fixed points in 2-dimensional maps. Then it is easy to understand that the parameter dependence of the central NHIMs has an enormous influence on the parameter dependence of the whole dynamics. Therefore, we are interested in a good understanding of the typical development scenarios of NHIMs under parameter changes. In autonomous Hamiltonian systems (the class of systems we deal with in the present article), the most important parameter is in many cases the total energy $E$, i.e. the conserved numerical value of the generating function $H$ of the dynamics.
@@
Hamiltonian systems have an effective potential; for the general recipe to construct it for a given Hamiltonian function, see [2, 3, 4]. And usually, the most important NHIMs of codimension 2 sit over the index-1 saddle points of this effective potential. Then these NHIMs act as transition states over these saddles and direct the flow over these saddles [5, 6, 7]. We can imagine that from parts of the stable and unstable manifolds of the NHIMs, we glue together tubes (cylinders) which guide the flow over the saddle of the effective potential. This property demonstrates the importance of the NHIMs in the generation of the global properties of the dynamics.
@@
For Hamiltonian systems with three degrees of freedom (3-dof), we are at the moment far away from a complete understanding of the typical development scenarios of codimension-2 NHIMs. We only have hard mathematical persistence theorems which state that NHIMs survive perturbations as long as the normal instability remains larger than the tangential instability, for various versions of proofs see [8], [9], chapter 3 in [1], and chapter 3 in [10]. Many numerical examples indicate that NHIMs really decay in regions where the tangential instability surpasses the normal instability. In the present situation, it is worth studying many examples of development scenarios of NHIMs numerically to gain more experience with the typical and frequent development scenarios.
@@
There are some investigations of the loss of normal hyperbolicity of invariant surfaces [11, 12, 13, 14]. Unfortunately, these publications use 2-dof systems or use a simplification which is equivalent to the reduction to 2 essential degrees of freedom. And they are mainly concerned with the consequences of the loss of normal hyperbolicity to transition state theory and to the flow rates over the saddles. They are not concerned with the disintegration of the NHIM surface into (partially lower-dimensional) invariant fragments, a topic into which we will delve in the present article. For some further observations of qualitative changes of NHIMs under parameter changes, see [15, 16, 17, 18].
@@
NHIMs of dimension 2 or more in a map have an internal dynamics which can be displayed graphically by plotting the Poincaré map restricted to the NHIM itself, which we will call the restricted map for short. For an early example of the use of such a restricted map, see Fig.1 in [11]. It is the same basic idea as the restriction of the Poincaré map to the centre manifold of the saddle point, see for example Fig.6 in [19]. For maps with 4 dimensions in total, this restricted map for NHIMs of codimension 2 is itself 2-dimensional and can be plotted conveniently. This restricted map, as a function of the parameters of the system, is an excellent way to display graphically the development scenario of the NHIMs. The numerical method to follow the NHIMs under parameter changes and to display the restricted map have been explained in detail in [20].
@@
In various examples, we have observed that in systems with several fundamental NHIMs the development scenarios of these NHIMs are coordinated and important qualitative changes of one particular NHIM occur in the same small parameter interval in which also some other NHIMs suffer important changes. One of these examples has been a simple model for the motion of a test particle in the effective potential of a rotating symmetric system of two spherical dwarf galaxies [21]. This system has three collinear Lagrange points and codimension-2 NHIMs over these saddle points of the effective potential. The two outer NHIMs are symmetry-related, whereas the middle one is qualitatively different from the outer ones. Nevertheless, the development scenario of the middle NHIM is coordinated with the development scenario of the two outer NHIMs. Now the reader can ask whether the discrete left-right symmetry of the complete system is responsible for this coordination of the development scenarios. Motivated by this question, we investigate in the present article the asymmetric generalisation of the model from [21]. In the present article, we set the masses of the two spherical dwarf galaxies to different values in contrast to [21], where these two masses have been equal.
@@
Another example of the occurrence of coordination effects in the development scenario of NHIMs has been observed in [22]. This system is the motion of a test particle in a rotating double-barred galaxy. The model has four codimension-2 NHIMs over index-1 saddles of the effective potential. These four NHIMs fall into two nonequivalent groups, where the two NHIMs within each group are symmetry related. Nevertheless, we found a coordination of the scenarios of all four NHIMs, i.e. also between the nonequivalent groups.
@@
In the treatment of these previous systems, we did not observe any important role of the transient outer parts of partially decayed NHIMs. In the meantime, two interesting cases of such transient effects have been observed, see [23] and [24]. Interestingly, in the asymmetric case of the two dwarf galaxies, the topic of the present article, the transient outer parts of the decaying NHIMs influence the coordination of the development and decay scenario and thereby of the creation and development of chaos. This is a new event, not noticed in previous studies. We consider it a remarkable new result of the investigation presented here. For general information on transients and in particular on transient chaos, see the textbook [25].
@@
In section 2, we introduce the asymmetric model and explain the important extremal points of its effective potential. Section 3 displays graphically the development scenario of the 3 NHIMs and of the most important periodic orbits related to these NHIMs. In section 4, we have a closer look at the creation of chaos and at heteroclinic connections between NHIMs, including the role of transient effects. In section 5, we use the delay time as an indicator function to obtain more information about the phase space structure, and in section 6 we draw conclusions and give final remarks.
@@ 2 The model
Here $c$ is a size measure for the galaxies, which we take equal for the two galaxies. The two galaxies rotate around each other and when $R$ is large compared to $c$ then we can approximate the rotational angular velocity $\Omega$ by the Kepler value
@@
In these units, the parameters of the system used in the following are $c=0.25$, $R=6$, $M_{1}=0.06$ and $M_{2}=0.14$. Note that accordingly $M_{t}=0.2$, exactly the same value as the one used for the symmetric case in [21]. Also, the values of $c$ and $R$ are chosen the same as in the symmetric case. We hope that this makes a comparison between the symmetric and asymmetric cases easier. The momenta conjugate to the Cartesian position coordinates are $p_{x},p_{y},p_{z}$. In these coordinates, the generating function of the dynamics (the Hamiltonian function) is given as
@@
Here $L_{z}=xp_{y}-yp_{x}$ is the $z$ component of the angular momentum. We describe the dynamics in a corotating system. Therefore, the numerical value of the function $H$ is not the energy of the test particle as measured by an observer in an inertial system. This energy, measured by the inertial observer, is not conserved because the potential seen by the inertial observer is explicitly time dependent. For the inertial observer, the conserved generating function from Eq.5 is the Jacobi conserved quantity. In the following, we use the letter $E$ for the numerical value of the function $H$, and call this number the energy, even though it is not the energy as measured by the inertial observer.
In later sections, a knowledge of some symmetries of this system will be useful. First, we notice that the effective potential is symmetric under $y\rightarrow-y$ and under $z\rightarrow-z$. If a trajectory starts with $z=0$ and $p_{z}=0$, then this trajectory will lie completely in the horizontal plane $z=0$ of the position space. Therefore, this horizontal plane is an invariant 2-dof subsystem. Because we describe the dynamics in a rotating frame, we do not find the usual time reversal symmetry. However, if we combine the transformation $t\rightarrow-t$ with a reflection in the plane $y=0$, then the orientation of rotation is also reversed and the Hamiltonian equations of motion remain the same. Therefore, the whole dynamics is invariant under the following transformation
@@
This transformation can be used to transform stable manifolds of invariant subsets into unstable manifolds of the corresponding $y$-reflected invariant subsets. This transformation will become useful in subsection 4b.
@@
Some contour lines of the effective potential from Eq.4 are plotted in the horizontal xyx-y plane in Fig.1. The points $L_{1}$, $L_{2}$, and $L_{3}$ are the collinear Lagrange points. In the 3-dimensional position space, they are index-1 saddles of $V_{eff}$ where the $x$ direction is the unstable direction. The 3 important NHIMs of the system sit over these saddle points. The Lagrange points $L_{4}$ and $L_{5}$ are index-2 saddle points of $V_{eff}$ and are of little importance in the following. The points $P_{1}$ and $P_{2}$ are relative minima of $V_{eff}$ and they lie close to the centres of the two galaxies because the distance $R$ is large compared to the size $c$. Compare this plot with Fig.3 in [21], which displays the corresponding symmetric case. In the following, we call $E_{j}$ the value of $V_{eff}$ in the Lagrange point $L_{j}$. And we use the names $E_{6}$ and $E_{7}$ for the values of $V_{eff}$ in the points $P_{1}$ and $P_{2}$ respectively. The numerical values of $E_{j}$ for our choice of the masses are: $E_{1}=-0.06511$..., $E_{2}=-0.05481$..., $E_{3}=-0.05921$..., $E_{4}=E_{5}=-0.04647$..., $E_{6}=-0.2714$... and $E_{7}=-0.5714$....
@@
Figure 1: The effective potential in the horizontal plane. The green, blue and red curves are the equipotential lines to the energies $E_{1}$, $E_{2}$ and $E_{3}$, respectively. Some additional equipotential curves are drawn as grey lines. The black dots mark the extremal points of the potential.
@@
In Fig.1, the equipotential curves to the energies $E_{1}$, $E_{2}$ and $E_{3}$ are given by the green, blue and red curves, respectively. The equipotential curves for some further energy values are plotted as grey curves. The extremal points of the potential are marked by black dots and labelled by their names given in the last paragraph.
@@
Finally, we give the motivation for our choice of the mass ratio between the two galaxies. As already mentioned, we want to keep $M_{t}$ fixed at the value 0.2. And then, it is clear that we have chosen $M_{1}=M_{2}=0.1$ for the symmetric case in [21]. And it is also obvious that in the symmetric case we had $E_{2}=E_{3}$ and also $E_{6}=E_{7}$, see Fig.3 in [21]. In the present article, we want to have the system strongly asymmetric and all three $E_{j}$ values of the collinear Lagrange points as distinct as possible. So we have chosen the values of the masses such that the distance between $E_{1}$ and $E_{3}$ is approximately equal to the distance between $E_{3}$ and $E_{2}$. Now the reader might ask: Why not use $M_{2}$ close to 0.2 and correspondingly $M_{1}=M_{t}-M_{2}$ close to zero? Then the ratio between $M_{2}$ and $M_{1}$ would be enormous. The problem is that in such a case, $E_{3}$ and $E_{1}$ would be very close together, and the saddle point $L_{3}$ would have little importance for the dynamics.
3 Numerical results for the development and decay scenario of the NHIMs
3.1 Development scenario of the most important periodic orbits
@@
From now on, we use the conserved numerical value $E$ of the Hamiltonian as the perturbation parameter of the system and study the development scenario of the dynamics as a function of $E$. Out of the index-1 saddles and out of the relative minima of $V_{eff}$ grow the most important periodic orbits of the system, which are partly also the most important periodic orbits within the NHIMs. Therefore, a detailed knowledge of the development scenario of these periodic orbits helps a lot to understand the development scenario of the NHIMs and, in the end, the development scenario of the whole dynamics. The important properties of these periodic orbits are displayed in the bifurcation diagram plotted in Fig.2 in an energy interval which is relevant for the topic of the present article.
@@
Figure 2: Bifurcation diagram of the most important periodic orbits. Red curves represent horizontal orbits. Dark green curves represent vertical orbits, blue curves represent tangentially stable tilted loop orbits, and the magenta curve represents tangentially unstable tilted loop orbits. Pitchfork bifurcations are marked as open black squares, and the loss of normal hyperbolicity is marked by eight-pointed black stars.
@@
The horizontal axis of the plot gives the $x$ coordinate of the periodic orbit at the moment when it crosses the plane $y=0$ of the position space in negative orientation. The vertical axis of the plot gives the value $E$ of the generating function $H$.
@@
Out of each minimum of the potential grows three fundamental periodic orbits, two of them are horizontal, i.e. completely contained in the plane $z=0$ of the position space, and one of them goes mainly into the vertical direction. One horizontal orbit has a negative sense of rotation, and the other one has a positive orientation of rotation. The horizontal orbits with positive orientation are not important for the development scenario of the dynamics and are not discussed in the following, nor included in Fig.2 in order not to overload the plot. In contrast, all the periodic horizontal orbits with a negative orientation and with period one with respect to the negative intersection with the plane $y=0$ are central for the development scenario and are included in Fig.2 as the red lines. There is one red curve growing out of each of the potential minima. Note that each one of these two red curves runs through a symmetry-broken pitchfork bifurcation, whereby each one turns into 3 red curves for higher values of the energy. For the curve coming out of the minimum $P_{1}$, this happens for $x\approx-3,E\approx-0.063$, and for the one growing out of the minimum $P_{2}$, it happens for $x\approx 2.8,E\approx-0.074$. For the event of symmetry-broken pitchfork bifurcations, see section 20.3 and Fig. 20.3.2 in [26].
@@
The approximately vertical periodic orbits are represented in Fig.2 as the dark green curves. Again, the $x$ value of the orbit at the moment of the negative intersection with the plane $y=0$ is plotted as a function of $E$.
@@
Out of each one of the 3 collinear Lagrange points grows one horizontal Lyapunov orbit with negative orientation of rotation in the xyx-y plane, again represented in Fig.2 as red curves, and one approximately vertical Lyapunov orbit, again represented in the plot as dark green curves. In the following, we call lhj and lvj the horizontal Lyapunov orbit and the vertical Lyapunov orbit, respectively, coming out of the Lagrange point $L_{j}$. These Lyapunov orbits are the most important and central periodic orbits within the NHIMs created over the collinear Lagrange points. And these Lyapunov orbits are born at $E_{j}$ as normally hyperbolic and tangentially elliptic, where the words normally and tangentially refer to the NHIM surface.
@@
Pitchfork bifurcations play an important role in the scenario, and they are indicated by open black squares in the plot. In such a pitchfork bifurcation, each one of the horizontal Lyapunov orbits splits off a pair of tilted loop orbits, which go out of the horizontal plane. The two members of such a pair are transformed into each other by a reflection in $z$ and are represented by a single blue curve in the Fig.2. For higher energy, these tilted loop orbits disappear in saddle-centre bifurcations with pairs of further tilted loop orbits (also represented by blue curves) being created in pitchfork bifurcations of the horizontal orbits coming out of the potential minima. The tilted loop orbits of this type created over the Lagrange point $L_{j}$ will be called ltj in the following. The orbit lh1 runs through a further pitchfork bifurcation at $E\approx-0.045$, where it splits off a further pair of tilted loop orbits which are represented by the magenta curve in the plot. These tilted loop orbits will be called lu1. With increasing energy, these new tilted loop orbits lu1 turn vertical very rapidly and are absorbed by the vertical Lyapunov orbit lv1 in an inverse pitchfork bifurcation at $E\approx-0.043$.
The tangential stability properties of the periodic orbits contained within the NHIMs and their changes during all these bifurcations can be visualised best in the restricted Poincaré maps of the NHIMs. They will be explained in detail below in the subsections describing the development scenarios of the NHIMs.
@@
At still higher energy, the horizontal Lyapunov orbits coming out of the collinear Lagrange points are destroyed in saddle-centre bifurcations together with horizontal periodic orbits coming out of the potential minima or the horizontal orbits created in the broken pitchfork bifurcations suffered by the horizontal orbits coming out of the potential minima. Shortly before their destruction, the horizontal Lyapunov orbits lose their normal hyperbolcity; their loss of normal hyperbolicity is marked in the plot by a black eight-pointed star.

### v2

# The NHIM bifurcation scenario of a particle in an asymmetric binary system of dwarf galaxies

@@ Abstract
We study the bifurcation scenario of a three-degree-of-freedom Hamiltonian system, a model based on the Lagrange restricted 3-body problem: a test particle moving in the gravitational field of a pair of interacting dwarf galaxies. The phase space of this system has 3 fundamental normally hyperbolic invariant manifolds (NHIMs) and their invariant stable and unstable manifolds form homoclinic/heteroclinic tangles. As the perturbation parameter increases, the NHIMs begin to lose normal hyperbolicity and their constituent KAM tori break, creating transient chaotic dynamics around them. We also observe a certain kind of coordination between the bifurcation scenarios of these NHIMs. We analyse this phenomenon using Poincaré maps and the delay time function.
@@ 1 Introduction
In chaotic dynamical systems, geometrical structures in the phase space, such as unstable invariant subsets, determine the global dynamics and, in particular, the creation of chaos to a large extent. As the simplest example, we can think of the unstable fixed points in a 2-dimensional map whose stable and unstable manifolds build up the chaotic invariant set, which can be displayed pictorially as a horseshoe construction traced out by the intersecting stable and unstable manifolds of the most important unstable fixed points. In maps with a domain of higher dimension, the role played in 2-dimensional maps by unstable fixed points is taken over by normally hyperbolic invariant manifolds (standard abbreviation NHIMs) of codimension 2; for general information on NHIMs see [1]. These invariant subsets have stable and unstable manifolds of codimension 1, which can form walls, channels, and tubes in the phase space, thereby directing the global dynamics. This is the direct generalisation of the role of stable and unstable manifolds of hyperbolic fixed points in the phase space of 2-dimensional maps. It is then easy to understand that the parameter dependence of the main NHIMs greatly influences the parameter dependence of the entire dynamics. Therefore, we are interested in a good understanding of the typical bifurcation scenarios of NHIMs under parameter changes. In autonomous Hamiltonian systems, the most important parameter is, in many cases, the energy $E$, i.e., the conserved numerical value of the generating function $H$ of the dynamics.
@@
Hamiltonian systems have an effective potential; for the general recipe to construct it for a given Hamiltonian function, see [2, 3, 4]. And usually, the most important NHIMs of codimension-2 are associated with the index-1 saddle points of this effective potential. Then, these NHIMs act as transition states of the potential energy saddles and direct the flow through these saddles [5, 6, 7]. Due to the uniqueness of solutions of ODEs, in the phase space, two different trajectories cannot cross. Then segments of the stable/unstable manifolds form cylinders that direct the trajectories through the saddle effective potential. This property demonstrates the importance of the NHIMs for the global properties of the dynamics.
@@
For Hamiltonian systems with three degrees of freedom (3-dof), the understanding of the typical bifurcation scenarios of codimension-2 NHIMs is far from complete. There are only mathematical theorems which state that NHIMs are persistent under perturbations as long as the normal instability remains larger than the tangential instability. For various versions of proofs, see [8], [9], chapter 3 in [1], and chapter 3 in [10]. Recent computer-assisted algorithms to prove the existence and persistence of NHIMs are described in [11] and [12]. Many numerical examples indicate that NHIMs break in phase space regions where the tangential instability surpasses the normal instability. In the present situation, it is worth studying many examples of bifurcation scenarios of NHIMs numerically to gain more information about the typical and frequent bifurcation scenarios.
@@
In the literature, there are some investigations on the loss of normal hyperbolicity of invariant surfaces [13, 14, 15, 16]. These publications use 2-dof systems or a simplification equivalent to the reduction to 2-dof. These works are mainly concerned with the consequences of the loss of normal hyperbolicity to transition state theory and to the flow rates through the saddles. However, there is not much literature about the NHIMs breaking into (partially lower-dimensional) invariant fragments, a topic into which we will delve in the present article. For some further observations of qualitative changes of NHIMs under parameter changes, see [17, 18, 19, 20].
@@
NHIMs of dimension 2 or more in the phase space have an internal dynamics which can be displayed graphically by plotting the Poincaré map. For an early example of the use of such a Poincaré map, see Fig.1 in [13]. It is a similar idea to the Poincaré map for the centre manifold of the saddle point, see for example Fig.6 in [21]. For Poincaré maps with 4 dimensions in total, the Poincaré map for NHIMs of codimension 2 is itself 2-dimensional and can be plotted conveniently by projecting it on a canonical plane. This projected Poincaré map, as a function of the system parameters, is a simple way to display the bifurcation scenario of the NHIMs. The numerical method to follow the NHIMs under parameter changes and to display the projected Poincaré map has been explained in detail in [22] and in the appendix.
@@
In various examples, we have observed that in systems with several fundamental NHIMs, the bifurcation scenarios of these NHIMs are coordinated and important qualitative changes of one particular NHIM occur in the same small parameter interval in which also some other NHIMs experience important changes. One of these examples has been a simple model for the motion of a test particle in the effective potential of a rotating symmetric system of two spherical dwarf galaxies [23]. This system has three collinear Lagrange points and codimension-2 NHIMs associated with these saddle points of the effective potential. The two outer NHIMs are symmetry-related, whereas the middle one is qualitatively different from the outer ones. Nevertheless, the bifurcation scenario of the middle NHIM is coordinated with the bifurcation scenario of the two outer NHIMs. Now the reader can ask whether the discrete left-right symmetry of the complete system is responsible for this coordination of the bifurcation scenarios. Motivated by this question, we investigate in the present article the asymmetric generalisation of the model from [23]. In the present article, we set the masses of the two spherical dwarf galaxies to different values in contrast to [23], where these two masses are equal.
@@
Another example of the occurrence of coordination effects in the bifurcation scenario of NHIMs has been observed in [24]. This system is the motion of a test particle in a rotating double-barred galaxy. The model has four codimension-2 NHIMs associated with index-1 saddles of the effective potential. These four NHIMs fall into two non-equivalent groups, where the two NHIMs within each group are symmetry related. Nevertheless, we found a coordination of the scenarios of all four NHIMs, i.e. also between the non-equivalent groups.
@@
In the treatment of these previous systems, we did not observe any important role of the transient outer parts of partially broken NHIMs. In the meantime, two interesting cases of such transient effects have been observed, see [25] and [26]. Interestingly, in the asymmetric case of the two dwarf galaxies, the system studied in the present article, the transient outer parts of the broken NHIMs influence the coordination of the bifurcation scenario and thereby of the creation and development of chaos. This is a new event that was not previously noted in previous studies. We consider it a remarkable new result of the investigation presented here. For general information on transient dynamics and, in particular on transient chaos, see the textbook [27].
@@
In section 2, we introduce the asymmetric model and explain the important extremal points of its effective potential. Section 3 displays the bifurcation scenario of the 3 NHIMs and of the most important periodic orbits related to these NHIMs. In section 4, we have a closer look at the creation of chaos and at heteroclinic connections between NHIMs, including the role of transient effects in the dynamics. In section 5, we use the delay time as an indicator function to obtain more information about the phase space structure, and in section 6, we draw conclusions and give final remarks.
@@ 2 The model
Here, $c$ is a size measure for the galaxies, which we assume to be equal for the two galaxies. The two galaxies rotate around each other and when $R$ is large compared to $c$ then we can approximate the rotational angular velocity $\Omega$ by the Kepler value
@@
In these units, the parameters of the system used in the following are $c=0.25$, $R=6$, $m_{1}=0.06$, and $m_{2}=0.14$. Note that accordingly $m_{t}=0.2$, exactly the same value as the one used for the symmetric case in [23]. Also, the values of $c$ and $R$ are chosen the same as in the symmetric case. We hope that this makes a comparison between the symmetric and asymmetric cases easier. The momenta conjugate to the Cartesian position coordinates are $p_{x},p_{y},p_{z}$. In these coordinates, the generating function of the dynamics (the Hamiltonian function) is given as
@@
Here $L_{z}=xp_{y}-yp_{x}$ is the $z$ component of the angular momentum. We describe the dynamics in a corotating system. Therefore, the numerical value of the function $H$ is the Jacobi constant, denoted by the letter $E_{J}$.
In later sections, a knowledge of some symmetries of this system will be useful. First, we notice that the effective potential is symmetric under $y\rightarrow-y$ and under $z\rightarrow-z$. If a trajectory starts with $z=0$ and $p_{z}=0$, then this trajectory will lie completely in the horizontal plane $z=0$ of the position space. Therefore, this horizontal plane is an invariant 2-dof subsystem. Because we describe the dynamics in a rotating frame, we do not find the usual time reversal symmetry. However, if we combine the transformation $t\rightarrow-t$ with a reflection in the plane $y=0$, then the orientation of rotation is also reversed, and the Hamiltonian equations of motion remain the same. Therefore, the whole dynamics is invariant under the following transformation
@@
This transformation can be used to transform stable manifolds of invariant subsets into unstable manifolds of the corresponding $y$-reflected invariant subsets. This transformation will become useful in section 4.
@@
Some contour lines of the effective potential from Eq.4 are plotted in the horizontal xyx-y plane in Fig.1. The points $L_{1}$, $L_{2}$, and $L_{3}$ are the collinear Lagrange points. In the 3-dimensional position space, they are index-1 saddles of $V_{eff}$ where the $x$ direction is the unstable direction. The 3 important NHIMs of the system are associated with these saddle points. The Lagrange points $L_{4}$ and $L_{5}$ are index-2 saddle points of $V_{eff}$ and are of little importance in the following. The points $P_{1}$ and $P_{2}$ are relative minima of $V_{eff}$ and they lie close to the centres of the two galaxies because the distance $R$ is large compared to the size $c$. Compare this plot with Fig.3 in [23], which displays the corresponding symmetric case. In the following, we call $E_{Ji}$ the value of $V_{eff}$ in the Lagrange point $L_{i}$. And we use the names $E_{J6}$ and $E_{J7}$ for the values of $V_{eff}$ in the points $P_{1}$ and $P_{2}$ respectively. The numerical values of $E_{J1}$ for our choice of the masses are: $E_{J1}\approx-0.06511$, $E_{J2}\approx-0.05481$, $E_{J3}\approx-0.05921$, $E_{J4}=E_{J5}\approx-0.04647$, $E_{J6}\approx-0.2714$, and $E_{J7}\approx-0.5714$.
@@
Figure 1: The effective potential $V_{eff}$ in the horizontal plane. The green, blue and red curves are the equipotential lines to the energies $E_{J1}$, $E_{J2}$, and $E_{J3}$, respectively. Some additional equipotential curves are drawn as grey lines. The black dots mark the extremal points of the potential.
@@
In Fig.1, the equipotential curves to the energies $E_{J1}$, $E_{J2}$, and $E_{J3}$ are given by the green, blue, and red curves, respectively. The equipotential curves for some further Jacobi constant values are plotted as grey curves. The extremal points of the potential are marked by black dots and labelled by their names given in the last paragraph.
@@
Finally, we give the motivation for our choice of the mass ratio between the two galaxies. As already mentioned, we want to keep $m_{t}$ fixed at the value 0.2. And then, it is clear that we have chosen $m_{1}=m_{2}=0.1$ for the symmetric case in [23]. And it is also obvious that in the symmetric case we had $E_{J2}=E_{J3}$ and also $E_{J6}=E_{J7}$, see Fig.3 in [23]. In the present article, we want to have the system strongly asymmetric and all three $E_{J}$ values of the collinear Lagrange points as distinct as possible. So we have chosen the values of the masses such that the distance between $E_{J1}$ and $E_{J3}$ is approximately equal to the distance between $E_{J3}$ and $E_{J2}$. Now the reader might ask: Why not use $m_{2}$ close to 0.2 and correspondingly $m_{1}=m_{t}-m_{2}$ close to zero? Then the ratio between $m_{2}$ and $m_{1}$ would be enormous. The problem is that in such a case, $E_{J3}$ and $E_{J1}$ would be very close together, and the saddle point $L_{3}$ would have little importance for the dynamics.
3 Numerical results for the bifurcation scenario of the NHIMs
3.1 Bifurcation scenario of the most important periodic orbits
@@
From now on, we use the Jacobi constant $E_{J}$, the conserved numerical value of the Hamiltonian, as the perturbation parameter of the system and study the changes in the phase space as a function of $E_{J}$. Out of the index-1 saddles and out of the relative minima of $V_{eff}$ grow the most important periodic orbits of the system, which are partly also the most important periodic orbits within the NHIMs. Therefore, a detailed knowledge of the bifurcation scenario of these periodic orbits helps a lot to understand the bifurcation scenario of the NHIMs and, in the end, the changes of the whole dynamics. The important properties of these periodic orbits are displayed in the bifurcation diagram plotted in Fig.2 in an interval of the Jacobi constant which is relevant for the topic of the present article.
@@
Figure 2: Bifurcation diagram of the most important periodic orbits. On the panel (a), red curves represent horizontal orbits, dark green curves represent vertical orbits, blue curves represent tangentially stable tilted loop orbits, and the magenta curve represents tangentially unstable tilted loop orbits. Pitchfork bifurcations are marked as open black squares, and the loss of normal hyperbolicity is marked by eight-pointed black stars. On the panel (b), black curves represent periodic orbits which are unstable in the normal direction and stable in the tangential direction. Red curves represent periodic orbits which are unstable in the normal direction and also unstable in the tangential direction, where, however, the normal instability is larger than the tangential instability. Violet curves represent periodic orbits with larger tangential instability. And green curves represent periodic orbits outside of the NHIMs.
@@
The horizontal axis of the plot gives the $x$ coordinate of the periodic orbit at the moment when it crosses the plane $y=0$ of the position space in negative orientation. The vertical axis of the plot gives the value $E_{J}$ of the generating function $H$. This holds for both parts of the figure.
@@
Out of each minimum of the potential grow three fundamental periodic orbits, two of them are horizontal, i.e. completely contained in the plane $z=0$ of the position space, and one of them goes mainly into the vertical direction. One horizontal orbit has a negative sense of rotation, and the other one has a positive orientation of rotation. The horizontal orbits with positive orientation are not essential to understand the change of the dynamics studied in the present work and are not discussed in the following, nor included in Fig.2 in order not to overload the plot. In contrast, all the periodic horizontal orbits with a negative orientation and with period one with respect to the negative intersection with the plane $y=0$ are central for the change of the dynamics and are included in Fig.2(a) as the red lines. There is one red curve growing out of each potential minimum. Note that each one of these two red curves runs through a symmetry-broken pitchfork bifurcation, whereby each one turns into 3 red curves for higher values of the Jacobi constant. For the red curve coming out of the minimum $P_{1}$ this happens for $x\approx-3,E_{J}\approx-0.063$, and for the one growing out of the minimum $P_{2}$, it happens for $x\approx 2.8,E_{J}\approx-0.074$. For the event of symmetry-breaking pitchfork bifurcations, see section 20.3 and figure 20.3.2 in [28].
@@
The approximately vertical periodic orbits are represented in the Fig.2(a) as the dark green curves. Again, the $x$ value of the orbit at the moment of the negative intersection with the plane $y=0$ is plotted as a function of $E_{J}$.
@@
Out of each one of the 3 collinear Lagrange points grows one horizontal Lyapunov orbit with negative orientation of rotation in the xxyy plane, again represented in the Fig.2(a) as red curves, and one approximately vertical Lyapunov orbit, again represented in the Fig.2(a) as dark green curves. In the following, we call $lh_{i}$ and $lv_{i}$ the horizontal Lyapunov orbit and the vertical Lyapunov orbit, respectively, coming out of the Lagrange point $L_{i}$, $i=1,2,3$. These Lyapunov orbits are the most important and central periodic orbits within the NHIMs created from the collinear Lagrange points. And these Lyapunov orbits are created at $E_{J,i}$ as normally hyperbolic and tangentially elliptic, where the words normally and tangentially refer to the NHIM.
@@
Pitchfork bifurcations play an important role in the bifurcation scenario, and they are indicated by open black squares in the plot. In such a pitchfork bifurcation, each one of the horizontal Lyapunov orbits splits off a pair of tilted loop orbits, which go out of the horizontal plane. The two members of such a pair of tilted loop orbits are transformed into each other by a reflection in $z$ and are represented by a single blue curve in the Fig.2(a). For a higher Jacobi constant, these tilted loop orbits disappear in saddle-centre bifurcations with pairs of further tilted loop orbits ( also represented by blue curves ) in Fig.2(a), being created in pitchfork bifurcations of the horizontal orbits coming out of the potential minima. The tilted loop orbits of this type created from the Lagrange point $L_{i}$ will be called $lt_{i}$ in the following. The orbit $lh_{1}$ runs through a further pitchfork bifurcation at $E_{J}\approx-0.045$, where it splits off a further pair of tilted loop orbits which are represented by the magenta curve in Fig.2(a). These tilted loop orbits will be called $lu_{1}$. With increasing Jacobi constant, these new tilted loop orbits $lu_{1}$ turn vertical very rapidly and are absorbed by the vertical Lyapunov orbit $lv_{1}$ in an inverse pitchfork bifurcation at $E_{J}\approx-0.043$.
@@
At still higher Jacobi constant, the horizontal Lyapunov orbits coming out of the collinear Lagrange points are destroyed in saddle-centre bifurcations together with horizontal periodic orbits coming out of the potential minima or the horizontal orbits created in the broken pitchfork bifurcations suffered by the horizontal orbits coming out of the potential minima.
The stability properties of the periodic orbits contained within the NHIMs and their changes during all these bifurcations are essential for the bifurcation scenario of the NHIMs. Therefore, we illustrate them in Fig.2(b), which contains the same orbits as Fig.2(a), only plotted in other colours which indicate stability properties. Black represents periodic orbits within the NHIMs, with instability in the normal direction combined with stability in the tangential direction. Red represents orbits within the NHIMs with instability in the normal direction and also instability in the tangential direction, where, however, the normal instability is higher than the tangential instability. Violet indicates orbits originally growing out of the NHIMs, where the tangential instability has already become larger than the normal instability. Accordingly, the violet orbits no longer belong to the NHIMs. Finally, green indicates orbits outside of the NHIMs, they are not part of the following analysis. The transition from red to violet of a curve in Fig.2(b) indicates the loss of normal hyperbolicity of this orbit. In Fig.2(a), the loss of normal hyperbolicity of horizontal Lyapunov orbits is marked by 8-pointed black stars. The implications of this loss of normal hyperbolicities for the changes of the NHIMs will become evident below in the following subsections describing the bifurcation scenarios of the NHIMs.